# Using Large Language Models for Automatic Hint Ranking and Generation

### v1

@@ Abstract
The use of Large Language Models (LLMs) has increased significantly recently, with individuals frequently interacting with chatbots to receive answers to a wide range of questions. In an era where information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills. This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers. We first introduce a manually constructed hint dataset, WikiHint, which includes 5,000 hints created for 1,000 questions. We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts. We assess the effectiveness of the hints with human participants who try to answer questions with and without the aid of hints. Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings. Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves hint quality, and (c) encoder-based models perform better than decoder-based models in hint ranking.
1 Introduction
@@
Figure 1: Pipeline of WikiHint dataset generation. The numbers in arrows indicate the counts of output questions.
@@
In recent years, question answering (QA) systems have risen in importance, giving users the opportunity to ask arbitrary questions and gain answers to them Mavi et al. (2024); Karpukhin et al. (2020); Zhao et al. (2021); Abdel-Nabi et al. (2023). The rapid development of Large Language Models (LLMs) Gemini Team et al. (2023); OpenAI et al. (2023); Dubey et al. (2024) has without doubt contributed to this, as well as to many other Natural Language Problems (NLPs) Qin et al. (2024). While the benefits of using LLMs for current information access are clear [^1] While the current LLMs have still considerable weaknesses such as hallucinations, once can assume that, in the future, these problems will be mitigated to large extent, judging from the speed of the recent technology advancement., there are some worries in relation to their potential effect on human development. One such concern relates to the potential weakening of important cognitive skills of users like thinking, reasoning, and remembering due to the expected widespread use of automatic question answering technologies, in particular, ones backed by powerful AI solutions  Heersmink (2024). Users who will rely mainly on the solutions presented by AI’s, might also be discouraged to practice and improve their reasoning abilities Alfredo et al. (2024). For example, Darvishi et al. (2024) demonstrate that students are more likely to depend on AI assistance instead of learning from it. Moreover, Jošt et al. (2024) examine the impact of LLMs as an automated problem-solving technology on education and learning outcomes, demonstrating that such systems can negatively affect the development of learning skills. Furthermore, psychological studies confirm the importance of obtaining an answer independently, enhancing user self-confidence and encouraging further learning Bandura (2013). Letting users come up with the correct answers by themselves should then also contribute to the positive psychological effect, potentially increasing their self-confidence and motivation for learning Usher and Pajares (2006).
While there is no simple remedy to the aforementioned problem, we would like to promote an approach that involves humans in the answer finding process, by providing them with hints rather than direct answers. Hints are meant to serve as subtle clues to guide potential users towards a correct answer, without revealing the solution Hume et al. (1996). This should engage human’s cognitive abilities, potentially leading to forming new pathways in brains based on the received hints and already possessed knowledge. Automatically generating hints for user questions could be used as an alternative for those who prefer to find the answers themselves; much like long-distance walking is an alternative to using a car for people who wish to take additional effort for the benefit of staying healthy Panter et al. (2018).
@@
In this paper, we propose the first dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint which has been manually constructed and designed for both hint generation as well as evaluation. We next explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models. We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches. Finally, we assess the effectiveness of a novel evaluation method for hint ranking called HintRank and compare it with other automatic evaluation techniques. To sum up, we make the following contributions in this paper:
@@
  • We release the first manually created dataset called WikiHint for HG task containing 5,000 hints and 1,000 questions.
@@
  • We finetune and evaluate LLMs on WikiHint to assess the dataset quality and LLMs capabilities in hint generation and ranking.
@@
  • We present several novel observations, including the findings of a positive correlation between hint convergence and helpfulness, an inverse correlation between their length and helpfulness, and the superiority of the answer-aware approach over the answer-agnostic approach. In general, our research can contribute to fostering research in explainable AI for education Khosravi et al. (2022).
2 Related Work
@@
Automatic question answering (QA) Karpukhin et al. (2020); Zhao et al. (2021); Abdel-Nabi et al. (2023) and question generation (QG) Kurdi et al. (2020); Lu and Lu (2021); Zhang et al. (2021) have advanced quite much in the last years. These tasks have seen numerous different datasets Trischler et al. (2017); Rajpurkar et al. (2016); Zhang et al. (2020), and evaluation metrics Nema and Khapra (2018); Mavi et al. (2022) proposed. The research related to hint generation is however still scarce despite that hinting is a common mechanism used by humans for question answering, and that automatic hint generation could be regarded as the third missing task alongside the two established ones, QA and QG. The prior research focused mainly on generating hints for programming Price et al. (2019); Kochmar et al. (2022); Barnes and Stamper (2008); McBroom et al. (2021) and typically in the context of intelligent tutoring systems.
@@
Automatic hint generation for factoid questions was first addressed by Jatowt et al. (2023). However, the authors neither released a dataset nor utilized LLMs, focusing instead on hints generated from selected Wikidata [^3] https://www.wikidata.org/ predicates. Moreover, their work only considered an answer-aware setting and did not explore the hint ranking task. Subsequently, Mozafari et al. (2024) released the first synthetic dataset for hint generation (HG) called TriviaHG, which was automatically generated using LLMs. However, the automatic generation increases the likelihood of false information within the dataset, particularly due to the hallucination phenomenon of LLMs. The authors also introduced the first automatic evaluation method, called Convergence, for assessing hint quality. However, this method requires substantial computational resources as it relies on LLMs for evaluation. Interested readers can also refer to the recent survey of Jangra et al. (2024) who discuss various types of hints and challenges associated with hint generation and evaluation.
@@
To the best of our knowledge, WikiHint is the first manually curated dataset for the HG task, with questions and hints verified by humans. Our proposed automatic evaluation method is also the first for this task that does not rely on LLMs and is lightweight enough to be used locally.
3 WikiHint Dataset
@@
The absence of high-quality, verified hint datasets poses a significant challenge given the demanding data requirements of LLMs for their effective training. In this section, we outline the process for constructing WikiHint dataset. Figure 1 provides an overview of the pipeline of the dataset generation process, which we explore in detail in the following sections.
3.1 Question Sampling Module
@@
We incorporated AI-generated questions using ChatGPT OpenAI (2022) and questions from existing popular QA datasets such as SQuAD 2.0 Rajpurkar et al. (2018) and Natural Questions (NQ) Kwiatkowski et al. (2019). The following prompt was used for question generation:
@@
where WIKIPEDIA_LINK is the URL for the Wikipedia page corresponding to ANSWER. As mentioned above, we also selected questions from SQuAD 2.0 and NQ making sure that their answers had dedicated Wikipedia articles with sufficiently long content. Finally, we manually verified the correctness of all the selected questions discarding questions that are either too general or have incorrect answers. Table 6 in Appendix A compares the difficulty levels of questions based on their sources.
Figure 2: The HintRank method.
3.2 Hint Creation
@@
The crowdsourcing platform Amazon Mechanical Turk [^4] https://www.mturk.com/ was then used to distribute the hint creation task among multiple workers. The instructions shown to crowdworkers asked them to create five hints for a question and its associated Wikipedia article. After generating the hints, the workers were asked to rank them on a scale from 1 to 5, with 1 being the most helpful in finding correct answers and 5 being the least helpful. Figure 9 in Appendix A displays the annotators’ interface for hint generation while Figures 10, 11, and 12 provide additional views of the summarized instructions, detailed instructions, and the provided examples, respectively, to further assist the crowdworkers. Each data submission was subsequently reviewed manually for quality and was either approved or rejected based on their assessment. Table 14 in Appendix A shows the detailed criteria used for the selection process. The most common reasons for rejection were hints that directly revealed the answers (answer leakage) and hints that were single words instead of complete sentences. Among the 2,788 submissions reviewed, 1,788 were rejected and 1,000 were accepted. We prepared several attributes to be included for each question, answer, and hint, as shown in Tables 7, 8, and 9 in Appendix A. We discuss some of them in Section 4. Finally, we divided the questions and hints into train and test subsets. The train subset includes 4,500 hints for 900 questions, while the test subset contains 500 hints for 100 questions. Tables 10, 11, 12, and 13 in Appendix A show few examples of hints taken from the WikiHint dataset.

### v2

# WikiHint: A Human-Annotated Dataset for Hint Ranking and Generation

@@ Abstract.
The use of Large Language Models (LLMs) has increased significantly with users frequently asking questions to chatbots. In the time when information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills. This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers. We first introduce a manually constructed hint dataset, WikiHint, which is based on Wikipedia and includes 5,000 hints created for 1,000 questions. We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts. We assess the effectiveness of the hints with human participants who answer questions with and without the aid of hints. Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings. Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves quality of generated hints, and (c) encoder-based models perform better than decoder-based models in hint ranking.
Hint Dataset, Hint Evaluation, Hint Generation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
1. Introduction
Figure 1. Hints for a question from the WikiHint dataset, with their corresponding rankings (1 being the highest and 5 the lowest) which let users find the answer. Note that the opposite arrangement of the hints (from 1 to 5) would make the answer finding task easiest.
@@ 1. Introduction
Figure 2. Pipeline of WikiHint dataset generation. The numbers in the arrows indicate the counts of output questions. Qs and Hs denote the number of questions and hints, respectively.
@@
In recent years, question answering (QA) systems have risen in importance, offering users the opportunity to effortlessly receive answers to different types of questions (Mavi et al., 2024; Karpukhin et al., 2020; Zhao et al., 2021; Abdel-Nabi et al., 2023). The rapid development of Large Language Models (LLMs) (Gemini Team et al., 2023; OpenAI et al., 2023; Dubey et al., 2024) has mainly contributed to this, impacting also many other Natural Language tasks (NLPs) (Qin et al., 2024). While the benefits of using LLMs for current information access are quite evident, concerns have been raised in relation to their potential effect on human cognitive development. One such worry relates to the potential weakening of important cognitive skills of users like thinking, reasoning, and remembering due to the expected widespread use of the automatic question answering technologies, in particular, ones backed by powerful AI technologies  (Heersmink, 2024). Users who will rely mainly on the solutions presented by AI’s might also be discouraged to practice and improve their reasoning abilities (Alfredo et al., 2024). For example, Darvishi et al. (2024) demonstrate that students are more likely to depend on AI assistance instead of learning from it. Jošt et al. (2024) examine the impact of LLMs as an automated problem-solving technology on education and learning outcomes, demonstrating that such systems can negatively affect the development of learning skills. Furthermore, psychological studies confirm the importance of obtaining an answer independently, enhancing user self-confidence (Bandura, 2013). Letting users come up with the correct answers by themselves should then also contribute to the positive psychological effect, potentially increasing their self-confidence (Usher and Pajares, 2006).
While the problem of over-reliance on chatbots is complex and still not researched well, we wish to promote approaches that involve humans in the answer finding process. Providing users with hints rather than direct answers is one such approach. Hints are designed to offer subtle clues that guide users toward the correct answer without explicitly revealing it (Hume et al., 1996). Unlike the concept of framing (aka ”scattered conceptualization”) (Entman, 1993) which elevates particular pieces of information to promote a certain view or cause bias (Garrido-Muñoz et al., 2021; Färber et al., 2020), hints help humans to recall previously known information or guide them in the reasoning process. In this study, we define hints for factoid questions [^1] Factoid questions are those with answers in the form of entities such as locations, persons, etc. as clues that encourage users to think critically, reason, and draw on their knowledge to deduce the answer rather than having it explicitly provided. This approach aims to support and enhance cognitive skills such as critical thinking, reasoning, and memory. To ensure that hints are meaningful and widely accessible, we generate them based on external knowledge and well-established facts. For this purpose, we select Wikipedia as the primary source, as it is one of the most widely recognized and comprehensive repositories of factual information.
@@
In this paper, we propose the first manually curated dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint, which has been constructed and designed for hint generation, ranking and evaluation. Figure 1 shows a question from the WikiHint dataset, with its hints and the corresponding ranks reflecting the usefulness of each hint. We explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models. We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches. Finally, we assess the effectiveness of a novel evaluation method for hint ranking, called HintRank, and compare it with other automatic evaluation techniques. Ranking hints allows arranging them either from easiest to hardest which can be used for selecting the most useful hints for a given question, or from hardest to easiest (as shown in Fig. 1). The latter arrangement could find applications in quizzes, knowledge assessment scenarios, or in certain entertainment use cases (e.g., in games).
In summary, we make the following contributions in this paper:
@@
  • We release the first manually created dataset called WikiHint for the HG task containing 5,000 hints and 1,000 questions.
@@
  • We finetune and evaluate diverse LLMs on WikiHint to assess the dataset quality and LLMs’ capabilities in hint generation and ranking.
@@
  • We present several novel observations, including the findings of a positive correlation between hint convergence and helpfulness, an inverse correlation between hint length and helpfulness, and the superiority of the answer-aware approach over the answer-agnostic approach in hint generation.
2. Related Work
@@
Automatic question answering (QA) (Karpukhin et al., 2020; Zhao et al., 2021; Abdel-Nabi et al., 2023) and question generation (QG) (Kurdi et al., 2020; Lu and Lu, 2021; Zhang et al., 2021) have progressed a lot in the last years. These tasks have seen numerous different datasets (Trischler et al., 2017; Rajpurkar et al., 2016; Zhang et al., 2020), and evaluation metrics (Nema and Khapra, 2018; Mavi et al., 2022) proposed. The research related to hint generation is however still scarce despite the fact that hinting is a common mechanism used by humans for question answering, and that automatic hint generation could be regarded as the third missing task alongside QA and QG. The prior research however focused mainly on generating hints for programming (Price et al., 2019; Kochmar et al., 2022a; Barnes and Stamper, 2008; McBroom et al., 2021).
In the context of intelligent tutoring systems, automated, machine-learning-based methods have been shown to be effective in generating feedback and hints, reducing reliance on expert-crafted rules (Kochmar et al., 2020, 2022b). However, the effectiveness of LLMs in pedagogical tasks remains an open challenge. For instance, Tack and Piech (2022) investigated the teaching capabilities of LLMs such as Blender (Jiang et al., 2023) and GPT-3 (Brown et al., 2020) in educational dialogues and found that, while these models perform well in conversational uptake, they lack pedagogical effectiveness, particularly in providing helpful feedback. To address this limitation, several studies have proposed different approaches to improving LLM-generated feedback. For example, Kulshreshtha et al. (2022) leveraged a few-shot question generation strategy to create personalized feedback for mathematics problems, demonstrating improved student learning outcomes. Similarly, McNichols et al. (2023) explored the use of in-context learning and various LLMs for distractor generation in multiple-choice mathematics questions but found that the models still exhibit weaknesses, requiring further research for improvement. Additionally, Tonga et al. (2024) investigated the impact of different LLMs and prompting strategies on the quality of generated feedback and hints, highlighting that the effectiveness of LLM-generated assistance is highly dependent on the choice of model and the prompting approach used.
@@
Automatic hint generation for factoid questions was first addressed by Jatowt et al. (2023). However, the authors neither released a dataset nor utilized LLMs, focusing instead on hints generated from selected Wikidata [^3] https://www.wikidata.org/ predicates. Moreover, their work only considered an answer-aware setting in hint generation and did not explore the hint ranking task. Subsequently, Mozafari et al. (2024b) released the first synthetic dataset for hint generation (HG) called TriviaHG, which was automatically generated using LLMs. However, the automatic generation increases the likelihood of false information within the dataset, particularly due to the hallucination phenomenon of LLMs. The authors also introduced the first automatic evaluation method, called Convergence, for assessing hint quality. However, that method requires substantial computational resources as it relies on LLMs for evaluation. Mozafari et al. (2024a) also demonstrated that hints automatically generated from TriviaHG can serve as a context for QA systems, outperforming alternative approaches such as retrieval-based and generative methods in RAG. Interested readers can also refer to the recent survey of Jangra et al. (2024) who discuss various types of hints and challenges associated with hint generation and evaluation.
@@
To our knowledge, WikiHint is the first data set for the HG task with questions and their hints verified by humans. Our proposed automatic evaluation method for hint ranking is also the first for this task that does not rely on LLMs and is lightweight enough to be used locally.
3. WikiHint Dataset
@@
The absence of verified high-quality hint datasets poses a significant challenge given the demanding data requirements of LLMs for their effective training. In this section, we outline the process for constructing WikiHint dataset. Figure 2 provides an overview of the pipeline of the dataset generation process, which we explore in detail in the following sections.
Figure 3. The MTurk Worker interface for generating and ranking hints. \${question} represents the question presented to the worker, \${answer} is the corresponding answer, and \${link} provides the link to the relevant Wikipedia page.
3.1. Question Sampling Module
We first select questions that meet our criteria for hint generation.
3.1.1. Question Selection
@@
We incorporated 4,000 questions, including randomly selected questions from popular QA datasets such as SQuAD 2.0 (Rajpurkar et al., 2018) and Natural Questions (NQ) (Kwiatkowski et al., 2019), as well as AI-generated questions using ChatGPT (OpenAI, 2022). The following prompt was used for question generation with ChatGPT:
@@
Figure 4. The instructions for the hint generation and ranking tasks on the Amazon MTurk platform.
Table 1. A detailed description of attributes of a question.
Attribute Description
question The content of the question.
major The major category of the question.
minor The specific sub-category of the question.
entity Content of an entity.
ent_type Type of entity (e.g., GPE, PERSON).
start_index Start index of the entity in the question.
end_index End index of the entity in the question.
wikipedia_page_title The title of the Wikipedia page corresponding to the entity in question.
readability The readability score.
familiarity The familiarity score.
difficulty The difficulty level of the question.
@@
where WIKIPEDIA_LINK is the URL of the Wikipedia page corresponding to ANSWER. As mentioned above, we also selected questions from SQuAD 2.0 and NQ making sure that their answers had dedicated Wikipedia articles with sufficiently long content. After this filtering process, 3,247 questions remained for the next steps.
3.1.2. Question Verification
We manually verified the correctness of all the selected questions, discarding those that were too general or contained incorrect answers. Additionally, we ensured that all questions included a question mark, adding it where necessary. We also reviewed all Wikipedia links and corrected any character errors in the URLs. After this filtering process, 2,788 questions remained for the next steps.
3.2. Hint Generation Module
In this module, we generate hints and rank them.
3.2.1. Amazon MTurk
@@
The crowdsourcing platform Amazon Mechanical Turk [^4] https://www.mturk.com/ was then used to distribute the hint creation task among multiple workers. The instructions shown to crowdworkers asked them to create five hints for a question and its associated Wikipedia article. After generating the hints, the same workers were asked to rank them on a scale from 1 to 5, with 1 being the most helpful in finding correct answers and 5 being the least helpful. Figure 3 displays the annotators’ interface for hint generation while Figure 4 provides detailed instructions to further assist the crowdworkers.
Table 2. A detailed description of attributes of an answer.
Attribute Description
answer The actual answer.
entity Content of an entity as identified within the answer.
ent_type Type of entity.
start_index Start index of the entity in the answer.
end_index End index of the entity in the answer.
wikipedia_page_title The title of the Wikipedia page corresponding to the entity in the answer.
familiarity The familiarity score.
difficulty The difficulty level of the answer.
Table 3. A detailed description of attributes of a hint.
Attribute Description
hint Hint provided for the question.
source URL source of the hint.
entity Entities mentioned in the hint.
wikipedia_page_title The title of the Wikipedia page corresponding to the entities in the hint.
relevance The relevance score.
readability The readability score.
convergence The convergence score.
familiarity The familiarity score.
answer_leakage The answer leakage score.
rank Priority or helpfulness rating of the hint.
3.2.2. Hint Verification
Each data submission was then manually reviewed by three independent judges for quality and either approved or rejected based on their evaluation. The following section outlines the detailed criteria used in the selection process.
  • The hint must not explicitly include the exact answer.
  • The hint must be a sentence.
  • The hint must be specific, not generic.
  • The hint must be from the corresponding Wikipedia page.
  • The hint must have a unique rank.
The most common reasons for rejection were hints that directly revealed the answers (answer leakage) and hints that were single words instead of complete sentences. Among the 2,788 submissions reviewed, 1,788 were rejected and 1,000 were accepted.
3.2.3. Dataset Attributes
We prepared several attributes to include for each question, answer, and hint, as shown in Tables 1, 2, and 3, respectively. We discuss some of them in Section 4.
Two key attributes of a question are major and minor, which indicate the question type. To extract these attributes, we employ question classifiers, which categorize natural language questions into predefined classes based on their intended meaning. For example, the question Who is the director of the Godfather movie? falls into the PERSON:INDIVIDUAL class, while What is the capital of Austria? belongs to the LOCATION:CITY class. To develop an effective question classifier, we fine-tune the RoBERTa (Liu et al., 2019a) model using the TREC Question Classification dataset (Li and Roth, 2002), resulting in a model called QT Detector. This model achieves an accuracy of $92.8\%$ on the TREC Question Classification dataset [^5] The accuracy of the QT Detector on the TREC Question Classification dataset is $92.8\%$. Using QT Detector, questions are categorized into hierarchical classes with varying levels of granularity, encompassing 5 coarse-grained classes and 50 fine-grained classes.
Finally, we divide the questions and hints into train and test subsets. The train subset includes 4,500 hints for 900 questions, while the test subset contains 500 hints for 100 questions.
Figure 5. The HintRank method. First, the inputs are concatenated, and special tokens are added to format them for the BERT model. Finally, the BERT model determines which hint is more helpful.
Figure 6. The distribution of the Train and Test subsets.
4. Evaluation Approaches
In this section, we first discuss the evaluation metrics we use to assess the quality of hints, followed by a detailed introduction to HintRank, our proposed automatic ranking method.
4.1. Evaluation Metrics
Mozafari et al. (2024b) introduced five evaluation metrics for assessing hint quality: Relevance, Readability, Convergence, Familiarity, and Answer Leakage. While the authors did not propose automatic methods for Relevance, Readability, and Answer Leakage, in this paper, we employ automated approaches for these metrics. In the following, we provide a detailed description of all metrics.
Figure 7. Question difficulty based on the question types.
Relevance
This metric assesses the semantic connection between the hint and the question, ensuring that the hint remains meaningful and contextually appropriate. Hints can be considered a form of an answer since they provide explanations of the correct answer of the question. Based on this, one can evaluate the Relevance of a hint to its question as an Answer Relevance task (Es et al., 2024) - the task where the goal is to assess how pertinent the provided answer is to the target question. To compute the answer relevance metric, we employ DeepEval framework [^6] https://docs.confident-ai.com/ treating the hint as a kind of answer.
Readability
This metric measures the ease of understanding a hint or question, ensuring clarity and accessibility. It is crucial for making the hints effective without confusing users. To evaluate Readability (Liu and Lee, 2023), we finetune a RoBERTa (Liu et al., 2019b) model as a classifier on the OneStopEnglish dataset (Vajjala and Lučić, 2018). The finetuned model categorizes sentences into three classes: Beginner (0), Intermediate (1), and Advanced (2), reflecting their level of reading difficulty [^7] The accuracy of the readability estimator model is 62.3%.
Convergence
This metric measures how effectively a hint can narrow down or eliminate potential answers to the given question. To evaluate hints based on Convergence, we adopt the approach proposed by Mozafari et al. (2024b). In this method, candidate answers are first generated for a question. Then, the metric determines how many candidate answers are supported by the target hint. The convergence score is then derived on the basis of the capability of the hint to eliminate many incorrect candidate answers. We utilize two core models for convergence evaluation: LLaMA-3.1-8B and LLaMA-3.1-70B (Dubey et al., 2024).
Table 4. Statistics of WikiHint dataset.